# A Search for Classical Subsystems in Quantum Worlds

### v2

@@ Abstract
Decoherence and einselection have been effective in explaining several features of an emergent classical world from an underlying quantum theory. However, the theory assumes a particular factorization of the global Hilbert space into constituent system and environment subsystems, as well as specially constructed Hamiltonians. In this work, we take a systematic approach to discover, given a fixed Hamiltonian, (potentially) several factorizations (or tensor product structures) of a global Hilbert space that admit a quasi-classical description of subsystems in the sense that certain states (the “pointer states”) are robust to entanglement. We show that every Hamiltonian admits a pointer basis in the factorization where the energy eigenvectors are separable. Furthermore, we implement an algorithm that allows us to discover a multitude of factorizations that admit pointer states and use it to explore these quasi-classical “realms” for both random and structured Hamiltonians. We also derive several analytical forms that the Hamiltonian may take in such factorizations, each with its unique set of features. Our approach has several implications: it enables us to derive the division into quasi-classical subsystems, demonstrates that decohering subsystems do not necessarily align with our classical notion of locality, and challenges ideas expressed by some authors that the propensity of a system to exhibit classical dynamics relies on minimizing the interaction between subsystems. From a quantum foundations perspective, these results lead to interesting ramifications for relative-state interpretations. From a quantum engineering perspective, these results may be useful in characterizing decoherence free subspaces and other passive error avoidance protocols.
@@ I Introduction
However, is this too provincial an approach, born out of our familiarity with certain macroscopic objects? One could instead take the reverse approach and, starting from the global Hamiltonian, ask if/what subsystems that system is composed of. This only entails choosing an energy spectrum and selecting associated eigenstates. Is it always possible, given an arbitrary Hamiltonian, to describe a system composed of various well-defined subsystems? Could more than one such composition coexist as separate interpretations of the same global picture? To start with, this change in perspective poses the question, what do we actually mean by a subsystem? That is, given a potential factorization of the total Hilbert space, what properties should that factorization satisfy to define good subsystems?
@@
Such questions have been explored by a number of authors using several different approaches [1, 2, 3, 4, 5, 6, 7, 8, 9]. Here, we emphasize that these questions are related to the concepts of environment-induced superselection (einselection) and pointer states. Pointer states are states of a subsystem that stay robust to entanglement with the rest of the physical world, namely, ‘the environment’. Often, but not always, these pointer states coincide with localized solutions to the Schrödinger equation and correspond to the type of entities that one usually associates with subsystems. A paradigmatic example would be the coherent states of a quantum harmonic oscillator interacting with a large environment [10, 11]. Einselection is the dynamical selection of these pointer states due to the environmental monitoring of the system [12, 13].
Pointer states have a special role in understanding decoherence due to several features (see e.g. Refs. [14, 15] for pioneering work). First, while the environment essentially performs a quantum non-demolition measurement on a system that is in a pointer state, for systems in superpositions of pointer states, the environment acts as a “which-path” monitor and leads to the damping of interference terms [13]. This explains why we do not see interference for classical systems interacting with a large environment. But more than just playing the passive role of a reservoir, the environment actively “acts as a witness” by redundantly storing information about the state of a system. Since individual observers only probe limited fragments of the environment, this selective redundancy of the pointer states is what leads to the emergence of an “objective” classical reality. [^1] Here “objective” reality is defined as the notion that multiple initially oblivious observers agree on the state of the system without affecting the others’ outcomes. This idea that the dynamics select certain states to be redundantly encoded in the environment also goes under the title of “Quantum Darwinism” [16, 17, 18] However, despite it’s success in explaining several features of the quantum-to-classical transition, the decoherence program is built on the assumption that there is a particular “splitting” of the world into a system and an environment such that there exist pointer states.
As pointed out by Zurek [19],
@@
It is this question that we set out to answer in this paper. We take a deductive approach: we ask, given that we observe a Universe that is partitioned into subsystems, into what factors can the Hilbert space be split such that there exist pointer states? To understand this relation between pointer states and subsystems operationally, we introduce an algorithm which, given a Hamiltonian, $H$, as an input, seeks to find a system-environment split and its associated pointer states. We utilize a variational algorithm with a cost function that can be used to identify product states of a system which, when evolved under $H$, remain robust to entanglement. This operational perspective helps crystallize our identification of a subsystem with the ideas of einselection.
@@
Analytical studies in conjunction with numerical work using our algorithm have enabled us to identify different classes of phenomena (i.e. different forms of interactions and different families of pointer states) that satisfy our definitions. We have found that any Hamiltonian can admit a certain class of pointer states and system-environment splits built in a straightforward way from the eigenstates of the Hamiltonian. This class is interesting in its own right, and serves as a useful reference point as we explore other ways our requirements can be realized. We systematically delineate these in Section IV. Some interesting special cases include solutions where a pointer state exists for some, but not all, environment states; or that a certain state stays robust to entanglement longer than the typical decoherence time but its corresponding orthogonal state does not share this behavior.
@@
We have found that one can identify different system-environment splits that coexist within a given global system. This observation may be relevant for engineering quantum technologies (for example in identifying all the available decoherence-free subspaces of specific quantum systems). We also comment on interesting foundational implications that come up in the cosmological context where there is no external “observer”.
@@
This paper is structured as follows. In Section II, we introduce the notion of subsystems and pointer states. In Section III, we introduce our algorithm for finding pointer states. In Section IV, we provide a detailed analysis of our numerical findings and further details of the algorithm. Section V explores the implications of our results for certain interpretational questions. (The reader who wishes to focus on these might try starting with Section V and follow references therein to other parts of this paper.) In Section VI we digest our results and discuss what conclusions can be drawn. Appendix A gives the details of a specific block-diagonal form for the Hamiltonian which we show is possible for any Hamiltonian given suitably chosen subsystems. Appendix B contains details regarding the implementation of the numerical algorithm. Appendix C gives details of the statistics that emerge from our numerical search algorithm. This work has a broad range of motivations, from understanding the emergence of fundamental laws of physics, to discovering technical results which may be relevant for quantum technologies. We gather these reflections in Section VI.
@@ II What is a subsystem?
For this work we consider the reverse process, starting with some basis $\{|k\rangle_{w}\}$ that spans ${w}$ and using Eq. (1) to define the tensor product structure (TPS). Operating in this way, we could alternatively insert
@@
in the left side of Eq. (1), where $B$ is some unitary on ${w}$. This scheme generates a different TPS (or factorization) determined by $B$. In what follows we will use the set of all unitaries on ${w}$ to scan through different TPS’s. In this work we consider only finite systems.
Operationally, the TPS shows up in the various operators one constructs to do physics, such as observables and parts of the Hamiltonian that act on, say, subsystem $\mathcal{A}$ or subsystem $\mathcal{B}$, exclusively. Note that operators that are “local” in the $\mathcal{H_{A}}\otimes\mathcal{H_{B}}$ factorization need not stay local from the point of view of a different factorization $\mathcal{H_{A^{\prime}}}\otimes\mathcal{H_{B^{\prime}}}$. And though the eigenvalues of the global Hamiltonian $H_{w}$ are not changed by the operation of $B$, the eigenstates of $H_{w}$ can look very different when expressed in different bases reflecting different TPS’s.
@@
When applied to the Hamiltonian, the transformation $BHB^{\dagger}$ rotates the eigenvectors but leaves the eigenvalues invariant.
Here, we explore a notion of subsystems associated with the processes of decoherence and einselection. We identify cases where the dynamics of the world (i.e. the system plus the environment) admit states in the system Hilbert space that stay robust to entanglement with the environment. Concretely, we ask: starting from only the total energy eigenvalues, can we arrive at a “preferred” factorization where the dynamics of the world (i.e. the system plus the environment) admit states in the system Hilbert space that stay robust to entanglement with the environment? For this work, we constrain our explorations to fixed sizes of the individual Hilbert space factors. Additionally, we wish to learn the properties of the global energy spectrum to which these factorizations owe their existence.
Some have argued that these preferred bipartite factorizations can be identified by their propensity to admit localized solutions [2]. While this certainly coincides with what we experience classically, this criterion is insufficient for identifying preferred factorizations though it is possible that these “localized” [^3] This strictly means that $\Delta x$, the width of the wavefunction, remains small solutions emerge as by-products of other underlying physical mechanisms in certain settings. Consider, for example, the Adapted Caldeira-Leggett (ACL) model of Ref. [20] where a quantum simple harmonic oscillator (SHO) interacts with a random environment so that the Hamiltonian takes the form,
Then, in the strong interaction limit ($\alpha_{i}\ll\beta$), the eigenstates of the position operator serve as states that remain robust to entanglement with the environment; in the weak interaction limit, where the self-Hamiltonian dominates, the energy eigenstates of the SHO remain robust to entanglement; finally, in the intermediate case, the coherent states (eigenstates of the annihilation operator) emerge as the pointer states [11]. Of these, the second case, barring perhaps the ground state wavefunction, is explicitly non-local even though the Hamiltonian takes on a form that coincides with the typical classical description of a weakly coupled system. Thus, the admission of local states must be a consequence, not the underlying cause, of the emergence of preferred factorizations of the Hilbert space.
The so-called “pointer states” – states in the system space that stay robust to entanglement with the environment – are natural candidates to search for in characterizing various factorizations of the Hilbert space (as in Eq. (4)). Then, in certain settings, these pointer states may coincide with localized solutions to the Schrödinger equation. This is true in the ACL model discussed above where pointer states exist in all three regimes of coupling strengths, but they are “local” only in the strong and intermediate regime.
It is also worth mentioning that decoherence and einselection do not embody all the elements of the emergence of classical from quantum. Some of these points are made for example in Refs. [21, 22, 23, 24]. Furthermore, several aspects of the emergence of classical from quantum depend on a multipartite split to become fully realized (this is illustrated for example in the Quantum Darwinism program [16]). Still, decoherence phenomena as realized in bipartite system-environment factorization is an essential ingredient for most of such discussions, and that is the focus of our work here.
III Algorithmic definition of subsystems
The search for preferred factorizations is, in principle, straightforward: wander in the space of unitaries $B$ each of which defines a new factorization for the Hilbert space; for each choice of $B$ iterate over the space of initial states $|\psi\rangle_{w}$; and for each choice of $|\psi\rangle_{w}$ evolve your system to late times and check if your reduced density matrix stably exhibits high values of purity. Thus, to identify whether an initial state $|\psi\rangle_{w}$ evolving under a Hamiltonian $H_{w}$ admits a well-defined subsystem, the question is whether we can identify some basis, mathematically denoted by the unitary rotation $B$ from the working basis, such that on rotating the composite system $\rho_{w}(t)$ by $B$ for all times $t\leq T$ the reduced state is pure. More concretely, we consider the linear entanglement entropy (henceforth simply referred to as the entropy),
[…]
of an initially pure world state
@@ III Algorithmic definition of subsystems
we say that it is robust to entanglement. The initial state of the system, i.e. $\operatorname{Tr}_{e}[\rho_{w}^{B}(0)]$, which is robust to entanglement for a $B$, $H$ and $|\psi\rangle_{w}$ such that Eq. (9) holds, is known as a pointer state. Any system that admits pointer states we will say is a subsystem [^4] We stress that our use of the term ‘subsystem’ includes the notion that it is a good subsystem - namely one that for certain initial states is robust to entanglement. and the dynamical process which leads to pointer states is known as einselection [12].
Should one wish to measure Eq. (9), the continuous integral over all times is impractical. Therefore, we suggest that weak robustness to entanglement can be more practically evaluated through the discretized metric
@@
where $\Delta T$ is a finite time interval and $T_{\rm train}$ is a maximum training time. Eq. (10) serves as a cost function that is to be minimized in our hunt for subsystems.
In order for the cost function to be faithful, the number of time-steps should be large (empirically, we find this to be $k\gtrsim 2d_{w}$ where $d_{w}\equiv\dim{(\mathcal{H}_{w}})$). This is because, for a small number of time steps, it is always possible to find specific states that minimize the cost at only the times on which the algorithm has been trained but that these states, unlike true pointer states, do not generalize to arbitrary times. This situation is explicitly depicted in Fig. 1. One can visualize this cost landscape as one that has many local minima for small values of $k$, but that many of these “evaporate” as $k$ increases, leaving behind solutions that generalize to times outside the training set. If the ansatz for the unitary $B$ is chosen with care and in correspondence to the structure of the Hamiltonian, it is plausible that much fewer training times can be used. Fully characterizing this algorithm is however not the focus of this work.
Furthermore, for the faithfulness of the cost function, the set of training times needs to contain points sufficiently beyond the characteristic decoherence time of the Hamiltonian. This is important since the reduced density matrix of a generic initial product state will remain pure for $t\lesssim t_{\rm dec}$ where the decoherence time, $t_{\rm dec}$, is determined by $\lambda_{\rm max}$, the largest eigenvalue of the Hamiltonian, i.e. $t_{\rm dec}\equiv 1/\lambda_{\rm max}$. This effect of varyig $T_{\rm train}$ on the generalizability of the solution can also be seen in Fig. 1.
In our investigations, we will consider only time-independent Hamiltonians and restrict ourselves to worlds consisting of two-state systems (qubits) so that $d_{w}\equiv\dim(\mathcal{H}_{w})=2^{n_{w}}$, where $n_{w}$ is the total number of qubits. This has the practical advantage that the Hamiltonian can be decomposed into a sum of Pauli strings which provides a convenient method of identifying interaction and self terms in the Hamiltonian in arbitrary factorizations. We also restrict our exploration to fixed sizes of the system and environment Hilbert space with dimensions $2^{n_{s}}$ and $2^{n_{e}}$ respectively.
We run the search algorithm multiple times, each time using a random starting guess for the factorization, $B$, and the initial state, $|\psi(0)\rangle_{w}$. Note that we always start with initial product states in the fiducial factorization. Since entanglement is factorization dependent, such a state will generally be entangled in another arbitrary factorization $B$. That is,
[…]
@@
Finally, we stress that it may not always be possible to achieve a vanishing cost (in fact, numerically this will essentially never occur), in which case we adopt the “predictability sieve” criterion [25, 11] that is a weaker definition of subsystems: factorizations that admit approximate pointer states in the system space that are weakly, though not entirely, robust to entanglement with the environment.
For details regarding the computational implementation of the search algorithm, we refer the reader to Appendix B.
Figure 2: Best late time linear entropy achieved by the different optimization methods for various Hamiltonians. The columns correspond to the Hamiltonians described in Sec. IV.2. The rows correspond to the different initial state optimizations from Sec. IV.1; the unitary $B$ is optimized in each case.

### v3

@@ Abstract
Decoherence and einselection explain many features of the emergence classical world from an underlying quantum theory. However, the theory assumes a particular factorization of the global Hilbert space into constituent system and environment subsystems. In this work, given a fixed Hamiltonian we show that there can exist several factorizations (or tensor product structures) of a global Hilbert space that admit a quasi-classical description of subsystems in the sense that certain states (the “pointer states”) are robust to entanglement. We derive several analytical forms that the Hamiltonian may take in such factorizations, each with its unique set of features. Our approach enables us to derive the division into quasi-classical subsystems and demonstrates that decohering subsystems do not necessarily align with our classical notion of locality. These results have interesting ramifications for the interpretation of quantum mechanics and, more practically, may be useful in algorithmically characterizing decoherence free subspaces.
@@ I Introduction
However, is this too provincial an approach, born out of our familiarity with certain macroscopic objects? One could instead take the reverse approach and, starting from the global Hamiltonian, ask if/what subsystems that system is composed of. This only entails choosing an energy spectrum and selecting associated eigenstates. Is it always possible, given an arbitrary Hamiltonian, to describe a system composed of various well-defined subsystems? Could more than one such composition coexist giving separate interpretations of the same global picture? To start with, this change in perspective poses the question, what do we actually mean by a subsystem? That is, given a potential factorization of the total Hilbert space, what properties should that factorization satisfy to define good subsystems?
@@
Such questions have been explored by a number of authors using several different approaches [deutsch1985quantum-interpretation-basis, zanardi2001virtual-subsystems, zanardi2004quantum-lloyd-lidar, piazza2010glimmers, tegmark2015consciousness, Cotler:2017abq, carroll2021quantum-singh, zanardi2022operational, Stoica:2024nhh, zurek2025decoherence]. Here, we emphasize that these questions are related to the concepts of environment-induced superselection (einselection) and pointer states. Pointer states are states of a subsystem that stay robust to entanglement with the rest of the physical world, namely, ‘the environment’. Often, but not always, these pointer states coincide with localized solutions to the Schrödinger equation and correspond to the types of entity that one usually associates with subsystems. Einselection is the dynamical selection of these pointer states due to the environmental monitoring of the system [zurek2003decoherence-and-the-quantum-origins, schlosshauer2019quantum-dec-review]. A paradigmatic example of this dynamical selection would be the coherent states of a quantum harmonic oscillator interacting with a large environment [paz1993reduction-zurek-habib, zurek1993coherent-states-Habib-Paz].
Pointer states have a special role in understanding decoherence due to several features (see e.g. Refs. [zurek1981pointer, joos1985emergence-zeh] for pioneering work and [schlosshauer2019quantum-dec-review] for a review). One of these key features is that the pointer states evolve predictably (i.e. their evolution in the reduced system space is described by some suitable master equation)  [zurek1993preferred-predictability-sieve]. Furthermore, while the environment essentially performs a quantum non-demolition measurement on a system that is in a pointer state, systems in superpositions of pointer states rapidly decohere for systems in superpositions of pointer states, the environment acts as a “which-path” monitor and leads to the damping of interference terms [schlosshauer2019quantum-dec-review].
However, despite it’s success in explaining several features of the quantum-to-classical transition, the decoherence program is built on the assumption that there is a particular “splitting” of the world into a system and an environment such that there exist pointer states.
As pointed out by Zurek [zurek1998decoherence-existential-interpretation-rough-guide],
@@
It is this question that we set out to answer here. We take a deductive approach and ask, given a Hamiltonian, into what factors can the Hilbert space be split such that there exist pointer states? To understand this relation between pointer states and subsystems operationally, we introduce an algorithm which, given a Hamiltonian, $H$, as an input, seeks to find a system-environment split and its associated pointer states. We utilize a variational algorithm with a cost function that can be used to identify product states of a system which, when evolved under $H$, remain robust to entanglement. This operational perspective helps us identify subsystems that exhibit einselection. Furthermore, we discover several analytical forms of Hamiltonian that support einselection and categorize our numerical searches into these different forms.
@@
These analytical studies in conjunction with the numerical work using our algorithm have enabled us to identify different classes of phenomena (i.e. different forms of system-environment interaction and different families of pointer states) that satisfy our definitions. We have found that any Hamiltonian can admit a certain class of pointer states and system-environment splits built in a straightforward way from the eigenstates of the Hamiltonian. This class is interesting in its own right, and serves as a useful reference point as we explore other ways our requirements can be realized. We systematically delineate these classes in Section IV. Some interesting special cases include solutions where a pointer state exists for some, but not all, environment states; or that a certain state stays robust to entanglement longer than the typical decoherence time but its corresponding orthogonal state does not share this behavior.
@@
We have found that one can identify different system-environment splits that coexist within a given global system. This observation may be relevant for engineering quantum technologies (for example in identifying all the available decoherence-free subspaces of specific quantum systems). We also comment on interesting foundational implications that arise in the cosmological context where there is no external “observer”.
It is important to note that decoherence and einselection do not embody all the elements of the emergence of classical from quantum. Some of these points are made for example in Refs. [albrecht1994inflation-squeezing, Anglin:1996bb, polarski1996semiclassicality-starobinsky-decoherence-wo-decoherence, strasberg2023everything]. Furthermore, several aspects of the emergence of classical from quantum depend on a multipartite split to become fully realized. For example, the Quantum Darwinism program [zurek2009quantum-darwinism] recognizes that the environment actively “acts as a witness” by redundantly storing information about the state of a system. Since individual observers only probe limited fragments of the environment, this selective redundancy of the pointer states is what leads to the emergence of an “objective” classical reality. [^1] Here “objective” reality is defined as the notion that multiple initially oblivious observers agree on the state of the system without affecting the others’ outcomes  [zurek2009quantum-darwinism, ollivier2004objective, ollivier2005environment]. Still, decoherence phenomena as realized in bipartite system-environment factorization is an essential ingredient for most of such discussions, and that is the focus of our work here. In the conclusions we discuss how our work could be extended to the multipartite case.
We also note that many of the topics we discuss here can be explored using the consistent histories formalism [Griffiths:1984rx, Omnes:1988ek, Omnes:1988em, Omnes:1988ej, Omnes:1988fv, Omnes:1992ag, Gell-Mann:1992wkv, Dowker:1994ac, Gell-Mann:2011rif] At the end of Sect. V we connect specific results from this paper with earlier work using consistent histories.
@@
This paper is structured as follows. In Section II, we introduce the analytical notions of subsystems and pointer states. In Section III, we introduce our algorithm for finding pointer states. In Section IV, we provide a detailed analysis of our numerical and analytical findings. Section V explores the implications of our results for certain interpretational questions. In Section VI we digest our results and discuss what conclusions can be drawn. Appendix A clarifies our usage of certain terminology. Appendix B gives the details of a specific block-diagonal form for the Hamiltonian which we show is possible for any Hamiltonian given suitably chosen subsystems. Appendix C contains details regarding the implementation of the numerical algorithm. Appendix D gives details of the statistics that emerge from our numerical search algorithm. This work has a broad range of motivations, from understanding the emergence of fundamental laws of physics, to discovering technical results which may be relevant for quantum technologies. We gather these reflections in Section VI.
@@ II What is a subsystem?
For this work we consider the reverse process - we start with some basis $\{|k\rangle_{w}\}$ that spans w{w} and use Eq. (1) to define the tensor product structure (TPS). Operating in this way, we could alternatively insert
@@
in the left side of Eq. (1), where $B$ is some unitary on w{w}. This scheme generates a different TPS (or factorization) determined by $B$. In what follows we will use the set of all unitaries on w{w} to scan through different TPS’s. We consider only finite systems and limit our explorations to bipartite factorizations.
Operationally, the TPS arises in the various operators one constructs to do physics, such as observables and parts of the Hamiltonian that act on, say, subsystem $\mathcal{A}$ or subsystem $\mathcal{B}$, exclusively. Note that operators that are “local” in the $\mathcal{H_{A}}\otimes\mathcal{H_{B}}$ factorization need not stay local from the point of view of a different factorization $\mathcal{H_{A^{\prime}}}\otimes\mathcal{H_{B^{\prime}}}$.
@@
$\mathcal{H_{A}}\otimes\mathcal{H_{B}}\xrightarrow{B}\mathcal{H_{A^{\prime}}}\otimes\mathcal{H_{B^{\prime}}}$ (4)
When applied to the Hamiltonian, the transformation $BHB^{\dagger}$ leaves the eigenvalues invariant, though the eigenstates of $H$ can look very different when expressed in different bases reflecting different factorizations.
At the level of the state, the change of TPS reflects a change in the entanglement structure between the subsystems described by a ket, $|\psi\rangle_{w}$. Since entanglement is factorization dependent, a product state, will generally be entangled in another arbitrary factorization. That is,
where, for the rest of this paper and in keeping with the decoherence literature, we use the notation $\sigma/s$ and $\epsilon/e$ to label the “system” and “environment” Hilbert spaces, respectively. In this notation, Eq. (4) becomes $\mathcal{H}_{\sigma}\otimes\mathcal{H}_{\epsilon}\xrightarrow{B}\mathcal{H}_{s}\otimes\mathcal{H}_{e}$ . The basis $|i\rangle_{s}$ is, in general, not related by a unitary transformation to $|k\rangle_{\sigma}$ (and likewise for the environment bases) even though clearly the global bases in each representation of ${\cal H}_{w}$ are related by the global unitary, BB. [^3] Intuitively, this is much like the case in $\mathbb{R}^{3}$, where a coordinate system $\{\hat{e}^{I}\}$ can be rotated by a transformation $M\in SO(3)$ to a basis $\{\hat{\theta}^{I}\}$ though individual basis elements in each coordinate basis will not generally be related by a lower dimensional isometry in $SO(2)$.
As a concrete example of this phenomena one can consider the Bell states which are entangled in the $s$ and $e$ factorization but, for example, could be non-entangled in the $\sigma$ and $\epsilon$ factorization:
$|\Phi_{+}\rangle_{w}=\frac{1}{\sqrt{2}}\left(|0\rangle_{s}\otimes|0\rangle_{s}+|1\rangle_{s}\otimes|1\rangle_{s}\right)\equiv|0\rangle_{\sigma}\otimes|0\rangle_{\epsilon}$ (6)
$|\Phi_{-}\rangle_{w}=\frac{1}{\sqrt{2}}\left(|0\rangle_{s}\otimes|1\rangle_{s}-|1\rangle_{s}\otimes|1\rangle_{s}\right)\equiv|0\rangle_{\sigma}\otimes|1\rangle_{\epsilon}$
$|\Psi_{+}\rangle_{w}=\frac{1}{\sqrt{2}}\left(|0\rangle_{s}\otimes|1\rangle_{s}+|1\rangle_{s}\otimes|0\rangle_{s}\right)\equiv|1\rangle_{\sigma}\otimes|0\rangle_{\epsilon}$
$|\Psi_{-}\rangle_{w}=\frac{1}{\sqrt{2}}\left(|0\rangle_{s}\otimes|1\rangle_{s}-|1\rangle_{s}\otimes|0\rangle_{s}\right)\equiv|1\rangle_{\sigma}\otimes|1\rangle_{\epsilon}\,.$
That is, here we have chosen the $\sigma$ and $\epsilon$ tensor product structure such that the Bell states are product states. Correspondingly, the unitary $B$ that transforms from the computational basis of $s$ and $e$, i.e., $\{|0\rangle_{s}\otimes 0\rangle_{e},|0\rangle_{s}\otimes 1\rangle_{e},|1\rangle_{s}\otimes 0\rangle_{e},|1\rangle_{s}\otimes 1\rangle_{e}\}$, to the Bell basis, $\{|\Phi_{+}\rangle_{se}=|0\rangle_{\sigma}\otimes|0\rangle_{\epsilon},|\Phi_{-}\rangle_{se}=|0\rangle_{\sigma}\otimes|1\rangle_{\epsilon},|\Psi_{+}\rangle_{se}=|1\rangle_{\sigma}\otimes|0\rangle_{\epsilon},|\Psi_{-}\rangle_{se}=|1\rangle_{\sigma}\otimes|1\rangle_{\epsilon}\}$ relates these two factorizations. However, different choices in $B$ would lead to different factorizations.
Here, we explore a notion of subsystems associated with the processes of decoherence and einselection. Concretely, we ask: starting from only the total energy eigenvalues, can we arrive at a “preferred” factorization where the dynamics of the world (i.e. the system plus the environment) admit states in the system Hilbert space that stay robust to entanglement with the environment?
Some have argued that these preferred bipartite factorizations can be identified by their propensity to admit localized solutions [piazza2010glimmers]. While this certainly coincides with what we experience classically, this criterion is insufficient for identifying preferred factorizations though it is possible that these “localized” [^4] This strictly means that $\Delta x$, the width of the wavefunction, remains small solutions emerge as by-products of other underlying physical mechanisms in certain settings.
Consider, for example, the Adapted Caldeira-Leggett (ACL) model of Ref. [albrecht2023adaptedcalmodel] where a quantum simple harmonic oscillator (SHO) interacts with a random environment so that the Hamiltonian takes the form,
Then, in the strong interaction limit ($\alpha_{i}\ll\beta$), the eigenstates of the position operator serve as states that remain robust to entanglement with the environment; in the weak interaction limit ($\beta\ll\alpha_{i}$), where the self-Hamiltonian dominates, the energy eigenstates of the SHO remain robust to entanglement; finally, in the intermediate case, the coherent states (eigenstates of the annihilation operator) emerge as the pointer states [zurek1993coherent-states-Habib-Paz]. Of these, the second case, barring perhaps the ground state wavefunction, is explicitly non-local even though the Hamiltonian takes on a form that coincides with the typical description of a weakly coupled system. Thus, although pointer states exist in all three regimes of coupling strengths, they are “local” only in the strong and intermediate regimes (see [Paz_1999] for a related discussion). The admission of local states must therefore be a consequence, not the underlying cause, of the emergence of preferred factorizations of the Hilbert space.
III An algorithmic search for subsystems
We now make precise how one can iterate over different factorizations of the Hilbert space to algorithmically discover the quasi-classical subsystems of a Hamiltonian.
III.1 The optimization problem
This search is, in principle, straightforward: wander in the space of unitaries, $B$, each of which defines a new factorization for the Hilbert space; for each choice of $B$ iterate over the space of initial states, $|\psi\rangle_{w}$; and for each choice of $|\psi\rangle_{w}$ evolve your system to late times and check if your reduced density matrix stably exhibits high values of purity. Thus, to identify whether an initial state, $|\psi\rangle_{w}$, evolving under a Hamiltonian, $H_{w}$, admits a well-defined subsystem, the question is whether we can identify some basis, mathematically denoted by the unitary rotation $B$ from the working basis, such that on rotating the composite system, $\rho_{w}(t)$, by $B$ for all times $t\leq T$, the reduced state is pure. More concretely, we consider the linear entanglement entropy of the subsystem
@@ III.1 The optimization problem
of an initially pure world state,
@@
we say that it is robust to entanglement. The initial state of the system, i.e. $\operatorname{Tr}_{e}[\rho_{w}^{B}(0)]$, which is robust to entanglement for a $B$, $H$, and $|\psi\rangle_{w}$ such that Eq. (11) holds, is identified as the pointer state. Any system that admits pointer states we will say is a subsystem [^5] We stress that our use of the term ‘subsystem’ includes the notion that it is a good subsystem - namely one that for certain initial states is robust to entanglement..
This, in principle, is the optimization problem to be solved. For practical reasons we use this discrete version of Eq. (11):
@@
where $\Delta t$ is a finite time interval and $T_{\rm train}$ is a maximum training time. Eq. (12) serves as the cost function that is to be minimized, by optimizing the initial state $|\psi(0)\rangle_{w}$ and the factorization $B$ for a given $H$, in our hunt for subsystems. We delineate different approaches to the initial conditions in the following subsection.
@@
It may not always be possible to achieve a vanishing cost (in fact, numerically this will essentially never occur), in which case we adopt the “predictability sieve” criterion [zurek1993preferred-predictability-sieve, zurek1993coherent-states-Habib-Paz] that is a weaker definition of subsystems: factorizations that admit approximate pointer states in the system space that are weakly, though not entirely, robust to entanglement with the environment.
We implement techniques from density matrix renormalization group theory to optimize the initial state and factorization (both parameterized by unitaries) the details of which can be found in Appendix C.1. However, more broadly, our general algorithmic approach, namely minimizing Eq. (12), is flexible and could be implemented in a number of different ways. Indeed, with the development of quantum hardware, the cost could eventually be measured directly on a quantum computer. Nonetheless, a reader interested in the implementation details of our optimization algorithm may refer to Appendix C.1.
In our investigations, we restrict ourselves to only bi-partite factorizations of finite-dimensional Hilbert spaces that are constructed from two-state systems (qubits) so that $d_{w}\equiv\dim(\mathcal{H}_{w})=2^{n_{w}}$, where $n_{w}$ is the total number of qubits. We also restrict our exploration to fixed sizes of the individual system and environment Hilbert space with dimensions $2^{n_{s}}$ and $2^{n_{e}}$ respectively.
III.2 Approaches to initial conditions
While we are interested in finding the optimal factorization, parameterized by the unitary $B$, our cost, Eq. (12), is a function of both $B$ and the initial state $|\psi(0)\rangle_{w}$. Thus, while we always optimize over $B$, we explore various levels of optimizing the initial state which may correspond to different physical situations. We outline each approach below in increasing order of complexity. As noted in Section I, this work has a range of motivations from constructing decoherence-free subspaces for quantum technologies to understanding the quantum-to-classical transition on cosmological scales. Thus, we also elaborate on how these different motivations are materialized in our choices for the initial conditions below.
No initial state optimization.
Here we fix our initial state to be a product state in the fiducial basis $|\psi(0)\rangle_{w}=|\xi\rangle_{\sigma}|\phi\rangle_{\epsilon}$ and optimize over just the factorization parameterized by $B$.
System state optimization.
In addition to optimizing over $B$, we also optimize over the initial state in the fiducial system Hilbert space, $\mathcal{H}_{\sigma}$, i.e. $|\psi(0)\rangle_{w}=A_{\sigma}\otimes 1_{\epsilon}(|\xi\rangle_{\sigma}|\phi\rangle_{\epsilon})$ where $A_{\sigma}\in U(d_{s})$ is to be optimized.
Global state optimization.
In this approach, in addition to optimizing over the factorization parameterized by $B$, we optimize the global initial state, i.e $|\psi(0)\rangle_{w}=A_{w}|\xi\rangle_{\sigma}|\phi\rangle_{\epsilon}$ where $A_{w}\in U(d_{w})$.
The first of the three initialization schemes is to address whether a set of dynamics, with a particular initial condition, can be viewed as corresponding to two approximately well-defined subsystems. The second case, where only the system state in the fiducial TPS can be varied, corresponds to the case encountered in the lab: the experimentalist can set the state of the system being studied, for example the spin of an electron, and not of the more disordered environment with which it interacts.
Our final optimization scheme is aimed at relieving ourselves of any prior notion of a working basis and, in cases where a notion of a system-environment factorization is not defined a priori, allows us to define the factorization in which a random initial system state decoheres. This last case especially pertains to cosmology [^6] As an illustration, there are two ambiguities that arise in studying the decoherence of curvature perturbations that are setup during the inflationary epoch. One is in deciding the cut-off scale for Fourier modes that form the system versus the environment. For example, some have defined the modes that leave the Hubble horizon (i.e. the ones that leave observable imprints on large-scale structure) as the “system” while modes that stay within the Hubble horizon as the “environment”, though this choice remains arbitrary [kiefer1998emergence, burgess2006decoherence-holman, martineau2007decoherence, sharman2007decoherence, burgess2023minimal-holman] (see also [polarski1996semiclassicality-starobinsky-decoherence-wo-decoherence] where it is argued that it is not necessary to specify an “environment” in order to explain the emergent classicality for inflationary perturbations which is due to quantum squeezing [albrecht1994inflation-squeezing] in what the authors refer to as “decoherence without decoherence”. The second ambiguity is in deciding the initial state in which the fluctuations originate. This is often taken to be the canonical “Bunch-Davies” vacuum [bunch1978quantum], but there are other choices (which, in some sense, may even be more “natural” and may have differing observable effects [Adil2023entanglement])., where there is no reason that a fiducial factorization should be privileged over another one.